Shirley’s mother is four times as old as Shirley is now. 10 years ago is Shirley’s mother was one more than three times Shirley’s age now.
Isnt this: 4x-10+1=3x??

Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

let x be Shirley's age now
:
Shirley's mother age now is 4x
:
3x +1 = 4x - 10
:

Answer by MathTherapy(10552)


Shirley’s mother is four times as old as Shirley is now. 10 years ago is Shirley’s mother was one more than three times Shirley’s age now.
Isnt this: 4x-10+1=3x??
No!! It's 4x - 10 = 3x + 1.